# Mapping Overlaps in Benchmarks through Perplexity in the Wild

### v2

@@ Abstract
We develop signatures of capacity familiarity to characterize large language model (LLM) benchmarks and their meaningful overlaps. Benchmark signatures probe the capacity required for benchmark performance. We formally define them as a set of salient tokens drawn from in-the-wild, naturally authored corpora, where LLM token perplexity, reflecting more or less pre-training exposure, becomes highly predictive of LLM benchmark performance. Through a large-scale meta-evaluation, we extract benchmark signatures via stepwise forward selection with linear regressions across 32 LLMs and 88 benchmarks spanning diverse knowledge, coding, logic, instruction following, math, language, reasoning, and cultural/world modeling. Our analysis situates signatures in relation to both the semantic similarity of benchmark questions and the correlation of model performance. While performance overlaps are universally high and semantic overlaps remain confined to a narrow mid-range, benchmark signatures prove highly informative in capturing variation, overlap, and divergence. We observe overlap in knowledge and reasoning subtasks, whereas multilingual and cultural benchmarks exhibit less similarity, even compared to cross-task overlap. Notably, performance-level results are strongly influenced by benchmark-orthogonal factors such as question format, highlighting limitations in LLM generalization, the conflation of performance with ability, and issues inherent in current mainstream benchmark agreement studies. Benchmark signatures, however, remain robust to such effects. Ultimately, we identify cross-functional overlaps across logic, math, language, instruction following, and world modeling, with coding emerging as the least overlapping domain. Together, these findings provide mechanistic insights into benchmark validity and LLM sensitivities, and sketch the underlying landscape of interconnected LLM capabilities.
@@ 1 Introduction
Why do in-the-wild corpora effectively encode benchmark characteristics? The abilities measured by benchmarks - commonsense, factual memory, scientific reasoning, programming skills, and more - do not emerge out of thin air. They stem from the diverse real-world text patterns encountered by the model during pretraining. In-the-wild corpora, consisting of large-scale, naturally authored, multi-domain text and code (news, forums, encyclopedias, textbooks and notes, papers, documentation, blogs, and repositories), are produced for human communication rather than adapted for benchmark design. They are rich in task-bearing structure (question–answer, problem–solution, claim–evidence, instruction–execution), redundancy (the same function expressed in many ways), and breadth. This breadth of distribution - likely unique to in-the-wild data - forms the “soil” from which such capabilities grow, and also the source from which benchmark questions are drawn. Even if a benchmark item never appears verbatim, its “function” recurs pervasively: unit-aware arithmetic in recipes (“double 1½ cups”), commonsense causality in narratives (“the glass shattered after being dropped”), claim $\rightarrow$ measurement $\rightarrow$ inference chains in scientific abstracts, code repair patterns in GitHub issues (“off-by-one in loop; fix bounds”), and even schema–query mappings (“customers with orders in last 30 days”). Focusing only on synthetic or benchmark-adjacent data risks capturing artifacts of test design. In-the-wild data, by contrast, mirrors the true distribution that gives rise to these abilities, making the overlap between pretraining exposure and benchmark competence not accidental but expected.
@@
Perplexity provides a useful lens for quantifying relationships between pretraining, skills, and benchmark performance. Low perplexity on a passage suggests that the model has seen similar linguistic and conceptual patterns during training and is familiar with the content. High perplexity, by contrast, indicates unfamiliarity and underrepresentation in pretraining. Thus, the distribution of perplexity values across large corpora serves as a fingerprint of the model’s training exposure and more or less acquired capacity. Importantly, because different benchmarks stress different capabilities, they map onto different perplexity distributions when probing across the same corpus. In other words, corpora encode benchmark signatures because benchmarks are not foreign entities imposed on the model after training, but rather structured samplings of capabilities that themselves emerge from the distribution of in-the-wild data. Perplexity serves as the bridge between corpus exposure and benchmark performance, making it possible to identify and characterize these signatures without requiring direct evaluation on the benchmark itself  [^2] More discussions about prior works see A.. We therefore leverage perplexity as the basis and covariate for salient token selection and signature formation  [^3] More details see Section 3 and Appendix B.. The following three levels in this work provide a holistic framework: semantics address task design, performance captures model behavior, and signatures reveal a fingerprint of model capacity. The overlap between benchmarks across each of these levels highlights the interconnected capacity space - an oft-discussed yet difficult-to-formalize concept and so represents a promising tool for evaluating benchmark validity. This rationale is illustrated in Figure 2.
@@
Figure 2: Overview of the rationale of how in-the-wild corpora implicitly encode the benchmark signature, knowledge exposure during pretraining (capacities), as well as benchmark performance.
@@
  • We develop a forward selection and regression-based pipeline to extract these signatures by mining and filtering token-level perplexity statistics from in-the-wild pretraining corpora.
@@ 3 Mining Benchmark Signatures from In-the-Wild Data
The overall process of mining signatures can be found in Algorithm 1 and its details can be found in Appendix 4. Let $d$ denote the number of “in-the-wild” tokens (in our case, tokens drawn from large-scale pretraining corpora [^4] Progressing from fine to coarse granularity, we have token-, chunk-, and document-level perplexities. We provide more experimental results of why the token-level operation is the best. Details are shown in Appendix B.), denoted as $\mathcal{T}=\{t_{1},\dots,t_{d}\}$, where $d$ typically scales to billions. For any benchmark $B_{j}$, our objective in extracting its benchmark signature is to isolate a subset of salient tokens $\mathcal{T}_{j}^{\prime}=\{t^{\prime}_{1},\dots,t^{\prime}_{d^{\prime}}\}\subset\mathcal{T}$ that are maximally informative in explaining variations in LLM performance. We formalize this as a regression problem: let $\mathbf{\hat{y}}_{j}:=(y_{1,j},\dots,y_{m,j})^{\top}\in\mathbb{R}^{m}$ denote the performance of $m$ language models on benchmark $B_{j}$. The covariate matrix $\mathbf{P}\in\mathbb{R}^{m\times d}$ contains token-level perplexities, where entry $\mathbf{P}_{ij}\equiv p_{ij}$ corresponds to the perplexity of token $t_{j}$ under LLM $M_{i}$. The challenge lies in the high-dimensional regime ($d\gg m$), where classical regression approaches are ill-posed. To make progress, we must uncover and exploit latent structural properties of the problem. In particular, we put forward a key assumption and a follow-up question:
@@ 3.1 Token-level Filtering with Perplexity Correlations
    Empirical Precedent in Data Selection: This screening methodology has demonstrated strong empirical success in the related domain of data selection for pre-training. Prior work has successfully used document-level perplexity correlations to filter large corpora, improving downstream model performance(Thrush et al., 2025; Shum et al., 2025). Their success provides compelling evidence for the practical utility of correlation screening as a robust heuristic for identifying informative signals in LLM-related data.
@@ 3.2 Refining Signatures with Forward Selection Regression
To guide this selection process, we use the Akaike Information Criterion (AIC; Bozdogan (1987)), which provides a principled trade-off between explanatory power and model size, mitigating the risk of overfitting. The process terminates when no additional token can improve the model’s AIC score by a meaningful amount. The complete two-stage process – combining the initial correlation screening to create a candidate set with the subsequent forward selection to derive the final signature – is formalized in Algorithm 4.
@@ 4.2 The Evaluation Bias from Model Family and Question Types is Resolved by the Signature
Grouping the result in Figure 4 panel b, we observe that the red areas are concentrated within the same benchmark family and question format as shown in Figure 1 right panel, where two red areas are exactly two benchmark families or question formats. We calculated pairwise correlations between benchmarks both within and across families and question formats. Since each family or format contains a highly diverse set of benchmarks - essentially covering everything - we would expect within-family/format overlaps to be quite low, showing little difference from cross-family/format overlaps. Consistent with this expectation, the signature-level analysis reveals statistically insignificant tiny differences based on the Mann–Whitney U test, yielding results around 0. This aligns with intuition, as the signature provides a good approximation of the true overlap and variation. In contrast, the performance-level analysis shows a large value of overlap (around 0.8) and a statistically significant increase in within-family/format overlap. Our results show that LLM performance may be more related to surface-level aspects of benchmarks such as question format, suggesting both that generalization and knowledge-propagation in LLMs are limited and that current evaluation may be underestimating peak performance because of conflation of performance and competence.
@@ 4.3 Signatures Inform Benchmark Design and LLM Capacity Space
As shown in Figure 1, we compare overlaps across design functions. Several patterns emerge. First, we observe significant overlaps that align with intuition. For example, math and logic correlate at 0.21, which is close to the average within-function overlap of 0.285 and far above the average cross-function overlap of 0.087. This makes sense: solving a math problem often requires logical reasoning, and vice versa. More broadly, logic, instruction following, language, math, and world modeling (largely cultural benchmarks) form a cluster of interconnected abilities. Coding appears far less entangled with other functions. Its low cross-function overlap suggests that coding benchmarks are comparatively “clean,” in the sense that success relies more specifically on coding competence and less on auxiliary abilities. This distinctiveness might arise because coding requires highly specialized pretraining corpora such as GitHub.
@@ Appendix A Related Literature
Signal Extraction from Pretraining Data: A growing body of work investigates how information extracted from pretraining corpora can inform data selection and model evaluation, even building benchmarks automatically. A central insight is that LLM losses on pretraining texts are often correlated with downstream benchmark performance, suggesting that simple loss-performance correlation coefficients can be effective signals for identifying high-quality pretraining data from in-the-wild corpus (Thrush et al., 2025; Hoffmann et al., 2022). Validation loss is thus frequently used as a proxy for model generalization (Kaplan et al., 2020; Hoffmann et al., 2022; Wei et al., 2022), and with more recent evidence showing that such correlations persist across architectures and training settings (Poli et al., 2023). One line of research focuses on efficient, low-cost methods for understanding and filtering signals, for instance lightweight approaches using surface-level heuristics (n-gram overlap (Xie et al., 2023) or semantic-level similarities (Everaert & Potts, 2023)), enabling scalable filtering of massive corpora. Thrush et al., (Thrush et al., 2025) proposed an orthogonal approach for data selection centered around estimates of perplexity-benchmark correlations. We build on these ideas to construct benchmark signatures by mining predictive tokens of LLM performance from large-scale pretraining corpora, in order to address challenges in meta-evaluation - the evaluation of LLM evaluations, e.g., how overlapping they are.
@@ Appendix B Comparisons between Token-, Chunk-, and Document-level Perplexity
From fine to coarse granularity, we consider token-, chunk-, and document-level perplexities. At the document level, we evaluate the model on an entire pretraining document and take the mean across all text chunks that fit within the model’s context window (part of the document). At the chunk level, we split documents into fixed-length windows (30 pieces, using spaces as separators) and compute perplexity as the average over all tokens within each window. At the token level—the finest granularity with the least inductive bias—we use token-wise perplexities from documents to capture the model’s intrinsic uncertainty. Concretely, we form a window by taking the target token with its up-to-30 preceding pieces (using spaces as separators) as context, then record only the last token’s perplexity as the feature. This ensures the token is conditioned on its preceding context rather than treated in isolation. As shown in Table 1, the token level exhibits the greatest standard deviation and interquartile range, as well as more pronounced extreme gaps in majority cases compared to the chunk and doc levels. This wider dispersion indicates that extreme values are more visible and significant at the token level, making it a natural choice for feature selection. Token-level signatures balance the strongest predictive power (both positive and negative relations) of highly informative tokens, and they exhibit high variance of predictive power, as captured by the deviations. By focusing on the token level, we are able to highlight more prominent signals, whereas aggregation at the chunk or document level tends to smooth out these extremes.

### v3

@@ Abstract
We develop signatures of capacity familiarity to characterize large language model (LLM) benchmarks and their meaningful overlaps. Benchmark signatures probe the capacity required for benchmark performance. We formally define them as a set of salient tokens drawn from in-the-wild, naturally authored corpora, where LLM token perplexity, reflecting more or less training exposure, becomes highly predictive of LLM benchmark performance. Through a large-scale meta-evaluation, we extract benchmark signatures via stepwise forward selection with linear regressions across 32 LLMs and 88 benchmarks spanning diverse knowledge, coding, logic, instruction following, math, language, reasoning, and cultural/world modeling. Our analysis situates signatures in relation to both the semantic similarity of benchmark questions and the correlation of model performance. While performance overlaps are universally high and semantic overlaps remain confined to a narrow mid-range, benchmark signatures prove highly informative in capturing variation, overlap, and divergence. We observe overlap in knowledge and reasoning subtasks, whereas multilingual and cultural benchmarks exhibit less similarity, even compared to cross-task overlap. Notably, performance-level results are strongly influenced by benchmark-orthogonal factors such as question format, highlighting limitations in LLM generalization, the conflation of performance with ability, and issues inherent in current mainstream benchmark agreement studies. Benchmark signatures, however, remain robust to such effects. Ultimately, we identify cross-functional overlaps across logic, math, language, instruction following, and world modeling, with coding emerging as the least overlapping domain. Together, these findings provide mechanistic insights into benchmark validity and LLM sensitivities, and sketch the underlying landscape of interconnected LLM capabilities.
@@ 1 Introduction
Why do in-the-wild corpora effectively encode benchmark characteristics? The abilities measured by benchmarks - commonsense, factual memory, scientific reasoning, programming skills, and more - do not emerge out of thin air. They stem from the diverse real-world text patterns encountered by the model during training. In-the-wild corpora, consisting of large-scale, naturally authored, multi-domain text and code (news, forums, encyclopedias, textbooks and notes, papers, documentation, blogs, and repositories), are produced for human communication rather than adapted for benchmark design. They are rich in task-bearing structure (question–answer, problem–solution, claim–evidence, instruction–execution), redundancy (the same function expressed in many ways), and breadth. This breadth of distribution - likely unique to in-the-wild data - forms the “soil” from which such capabilities grow, and also the source from which benchmark questions are drawn. Even if a benchmark item never appears verbatim, its “function” recurs pervasively: unit-aware arithmetic in recipes (“double 1½ cups”), commonsense causality in narratives (“the glass shattered after being dropped”), claim $\rightarrow$ measurement $\rightarrow$ inference chains in scientific abstracts, code repair patterns in GitHub issues (“off-by-one in loop; fix bounds”), and even schema–query mappings (“customers with orders in last 30 days”). Focusing only on synthetic or benchmark-adjacent data risks capturing artifacts of test design. In-the-wild data, by contrast, mirrors the true distribution that gives rise to these abilities, making the overlap between training exposure and benchmark competence not accidental but expected.
@@
Perplexity provides a useful lens for quantifying relationships between training, skills, and benchmark performance. Low perplexity on a passage suggests that the model has seen similar linguistic and conceptual patterns during training and is familiar with the content. High perplexity, by contrast, indicates unfamiliarity and underrepresentation. Thus, the distribution of perplexity values across large corpora serves as a fingerprint of the model’s training exposure and more or less acquired capacity. Importantly, because different benchmarks stress different capabilities, they map onto different perplexity distributions when probing across the same corpus. In other words, corpora encode benchmark signatures because benchmarks are not foreign entities imposed on the model after training, but rather structured samplings of capabilities that themselves emerge from the distribution of in-the-wild data. Perplexity serves as the bridge between corpus exposure and benchmark performance, making it possible to identify and characterize these signatures without requiring direct evaluation on the benchmark itself  [^2] More discussions about prior works see A.. We therefore leverage perplexity as the basis and covariate for salient token selection and signature formation  [^3] More details see Section 3 and Appendix B.. The following three levels in this work provide a holistic framework: semantics address task design, performance captures model behavior, and signatures reveal a fingerprint of model capacity. The overlap between benchmarks across each of these levels highlights the interconnected capacity space - an oft-discussed yet difficult-to-formalize concept and so represents a promising tool for evaluating benchmark validity. This rationale is illustrated in Figure 2.
@@
Figure 2: Overview of the rationale of how in-the-wild corpora implicitly encode the benchmark signature, knowledge exposure, as well as benchmark performance.
@@
  • We develop a forward selection and regression-based pipeline to extract these signatures by mining and filtering token-level perplexity statistics from large-scale in-the-wild corpora.
@@ 3 Mining Benchmark Signatures from In-the-Wild Data
The overall process of mining signatures can be found in Algorithm 1 and its details can be found in Appendix 4. Let $d$ denote the number of “in-the-wild” tokens (in our case, tokens drawn from large-scale in-the-wild corpora [^4] Progressing from fine to coarse granularity, we have token-, chunk-, and document-level perplexities. We provide more experimental results of why the token-level operation is the best. Details are shown in Appendix B.), denoted as $\mathcal{T}=\{t_{1},\dots,t_{d}\}$, where $d$ typically scales to billions. For any benchmark $B_{j}$, our objective in extracting its benchmark signature is to isolate a subset of salient tokens $\mathcal{T}_{j}^{\prime}=\{t^{\prime}_{1},\dots,t^{\prime}_{d^{\prime}}\}\subset\mathcal{T}$ that are maximally informative in explaining variations in LLM performance. We formalize this as a regression problem: let $\mathbf{\hat{y}}_{j}:=(y_{1,j},\dots,y_{m,j})^{\top}\in\mathbb{R}^{m}$ denote the performance of $m$ language models on benchmark $B_{j}$. The covariate matrix $\mathbf{P}\in\mathbb{R}^{m\times d}$ contains token-level perplexities, where entry $\mathbf{P}_{ij}\equiv p_{ij}$ corresponds to the perplexity of token $t_{j}$ under LLM $M_{i}$. The challenge lies in the high-dimensional regime ($d\gg m$), where classical regression approaches are ill-posed. To make progress, we must uncover and exploit latent structural properties of the problem. In particular, we put forward a key assumption and a follow-up question:
@@ 3.1 Token-level Filtering with Perplexity Correlations
    Empirical Precedent in Data Selection: This screening methodology has demonstrated strong empirical success in the related domain of data selection for training. Prior work has successfully used document-level perplexity correlations to filter large corpora, improving downstream model performance(Thrush et al., 2025; Shum et al., 2025). Their success provides compelling evidence for the practical utility of correlation screening as a robust heuristic for identifying informative signals in LLM-related data.
@@ 3.2 Refining Signatures with Forward Selection Regression
To guide this selection process, we use the Akaike Information Criterion (AIC; Bozdogan (1987)), which provides a principled trade-off between explanatory power and model size, mitigating the risk of overfitting. The process terminates when no additional token can improve the model’s AIC score by a meaningful amount. The complete two-stage process – combining the initial correlation screening to create a candidate set with the subsequent forward selection to derive the final signature – is formalized in Algorithm 4. To assess the robustness of our approach, we perform a leave-one-out cross-validation and compare it against a baseline that predicts performance using the mean of observed values. For each benchmark, we hold out one model, fit the regression on the remaining models, predict the held-out model outcomes, and repeat over all models. As a baseline, we predict the held-out value using the mean of the remaining models’ performances. The baseline produces an average MAE of 0.180, while our linear models achieve an average MAE of 0.075, demonstrating a substantial improvement over mean-based prediction.
@@ 4.2 The Evaluation Bias from Model Family and Question Types is Resolved by the Signature
Grouping the result in Figure 4 panel b, we observe that the red areas are concentrated within the same benchmark family and question format as shown in Figure 1 right panel, where two red areas are exactly two benchmark families or question formats. We calculated pairwise correlations between benchmarks both within and across families and question formats. Since each family or format contains a highly diverse set of benchmarks - essentially covering everything - we would expect within-family/format overlaps to be quite low, showing little difference from cross-family/format overlaps. Consistent with this expectation, the signature-level analysis reveals statistically insignificant tiny differences based on the Mann–Whitney U test, yielding results around 0. This aligns with intuition, as the signature provides a good approximation of the true overlap and variation. In contrast, the performance-level analysis shows a large value of overlap (around 0.8) and a statistically significant increase in within-family/format overlap. Our results show that LLM performance may be more related to surface-level aspects of benchmarks such as question format, suggesting both that generalization and knowledge-propagation in LLMs are limited and that current evaluation may be underestimating peak performance because of conflation of performance and competence. Using linear regression to obtain signature filters out the noise associated with the error term while preserving the underlying systematic relationships among benchmarks and performances.
@@ 4.3 Signatures Inform Benchmark Design and LLM Capacity Space
As shown in Figure 1, we compare overlaps across design functions. Several patterns emerge. First, we observe significant overlaps that align with intuition. For example, math and logic correlate at 0.21, which is close to the average within-function overlap of 0.285 and far above the average cross-function overlap of 0.087. This makes sense: solving a math problem often requires logical reasoning, and vice versa. More broadly, logic, instruction following, language, math, and world modeling (largely cultural benchmarks) form a cluster of interconnected abilities. Coding appears far less entangled with other functions. Its low cross-function overlap suggests that coding benchmarks are comparatively “clean,” in the sense that success relies more specifically on coding competence and less on auxiliary abilities. This distinctiveness might arise because coding requires highly specialized training corpora such as GitHub.
@@ Appendix A Related Literature
Signal Extraction from Training Data: A growing body of work investigates how information extracted from training corpora can inform data selection and model evaluation, even building benchmarks automatically. A central insight is that LLM losses on training texts are often correlated with downstream benchmark performance, suggesting that simple loss-performance correlation coefficients can be effective signals for identifying high-quality training data from in-the-wild corpus (Thrush et al., 2025; Hoffmann et al., 2022). Validation loss is thus frequently used as a proxy for model generalization (Kaplan et al., 2020; Hoffmann et al., 2022; Wei et al., 2022), and with more recent evidence showing that such correlations persist across architectures and training settings (Poli et al., 2023). One line of research focuses on efficient, low-cost methods for understanding and filtering signals, for instance lightweight approaches using surface-level heuristics (n-gram overlap (Xie et al., 2023) or semantic-level similarities (Everaert & Potts, 2023)), enabling scalable filtering of massive corpora. Thrush et al., (Thrush et al., 2025) proposed an orthogonal approach for data selection centered around estimates of perplexity-benchmark correlations. We build on these ideas to construct benchmark signatures by mining predictive tokens of LLM performance from large-scale in the wild corpora, in order to address challenges in meta-evaluation - the evaluation of LLM evaluations, e.g., how overlapping they are.
@@ Appendix B Comparisons between Token-, Chunk-, and Document-level Perplexity
From fine to coarse granularity, we consider token-, chunk-, and document-level perplexities. At the document level, we evaluate the model on an entire document and take the mean across all text chunks that fit within the model’s context window (part of the document). At the chunk level, we split documents into fixed-length windows (30 pieces, using spaces as separators) and compute perplexity as the average over all tokens within each window. At the token level—the finest granularity with the least inductive bias—we use token-wise perplexities from documents to capture the model’s intrinsic uncertainty. Concretely, we form a window by taking the target token with its up-to-30 preceding pieces (using spaces as separators) as context, then record only the last token’s perplexity as the feature. This ensures the token is conditioned on its preceding context rather than treated in isolation. As shown in Table 1, the token level exhibits the greatest standard deviation and interquartile range, as well as more pronounced extreme gaps in majority cases compared to the chunk and doc levels. This wider dispersion indicates that extreme values are more visible and significant at the token level, making it a natural choice for feature selection. Token-level signatures balance the strongest predictive power (both positive and negative relations) of highly informative tokens, and they exhibit high variance of predictive power, as captured by the deviations. By focusing on the token level, we are able to highlight more prominent signals, whereas aggregation at the chunk or document level tends to smooth out these extremes.
@@ E.2 Datasets
E.2.1 Models Used
Model (HF Repository) Citation
meta-llama/Llama-3.1-8B-Instruct Grattafiori et al. (2024)
meta-llama/Llama-3.2-1B-Instruct Grattafiori et al. (2024)
meta-llama/Llama-3.2-3B-Instruct Grattafiori et al. (2024)
google/gemma-3-4b-it Team (2025a)
google/gemma-3-12b-it Team (2025a)
google/gemma-3-27b-it Team (2025a)
mistralai/Mistral-7B-Instruct-v0.3 Jiang et al. (2023)
deepseek-ai/deepseek-llm-7b-chat DeepSeek-AI et al. (2024)
Qwen/Qwen3-0.6B Team (2025b)
Qwen/Qwen3-1.7B Team (2025b)
Qwen/Qwen3-4B Team (2025b)
Qwen/Qwen3-8B Team (2025b)
tiiuae/falcon-rw-1b Penedo et al. (2023)
EleutherAI/pythia-1b Biderman et al. (2023)
EleutherAI/pythia-6.9b-v0 Biderman et al. (2023)
EleutherAI/pythia-12b-deduped Biderman et al. (2023)
mosaicml/mpt-7b-instruct Team (2023)
microsoft/Phi-3-mini-4k-instruct Abdin et al. (2024a)
microsoft/Phi-4-mini-instruct Microsoft et al. (2025)
microsoft/Phi-4 Abdin et al. (2024b)
01-ai/Yi-1.5-9B-Chat AI et al. (2025)
01-ai/Yi-1.5-6B-Chat AI et al. (2025)
mistralai/Ministral-8B-Instruct-2410 Jiang et al. (2024)
openai-community/gpt2-medium Radford et al. (2019)
openai-community/gpt2-large Radford et al. (2019)
openai-community/gpt2-xl Radford et al. (2019)
zai-org/chatglm3-6b GLM et al. (2024)
zai-org/glm-4-9b-chat-hf GLM et al. (2024)
zai-org/codegeex4-all-9b Zheng et al. (2023)
allenai/OLMo-2-1124-13B-Instruct OLMo et al. (2024)
allenai/OLMo-2-1124-7B-Instruct OLMo et al. (2024)
allenai/OLMo-2-0425-1B-Instruct OLMo et al. (2024)
Table 4: Models used in this study. Multiple variants may share the same citation.